The following information will be used for the next 4 problems (18-21). I suggest you use Excel to do the calculations. A key hospital supplier, IVs Plus (IVP) located in Salina, KS, sells IV tubing and stands to hospitals and clinics. Sales have picked up ever since they introduced their newest Squeaky Clean IV stand, which eliminates all oils and germs left behind by users. Though IVP sells these stands all year long, they sell the most during the summer months, when end-of-fiscal year purchases are at a peak. The demand over the next 12 months is shown in the table below. Use the demand forecasts and determine the lowest cost production plan. Month Demand Forecast Month Demand Forecast January 133,067 July 251,630 February 155,026 August 249,630 March 168,200 September 200,312 April 173,890 October 160,830 May 202,759 November 145,266 June 260,842 December 128,900
Answer #1

Sum of total demand = 2230352

Begining inventory = 430000

Required production = 2230352 - 430000 =1800352= Less than 1900000

Number of employees required = Required production / (9600*12) =15.62 = 16

Employees hired = 0 , Employees fired = 18-16 = 2

Firing cost = $16000 * 2 = $32000, Hence between 30,001 and 40000